How many number of strings that can be formed by ordering the letters given:

(a) SCHOOL

(b) CLASSICS 


1
Expert's answer
2021-11-08T02:53:34-0500

Let us use the formula for permutations with repetitions.


(a) The number of strings that can be formed by ordering the letters of the string SCHOOL is equal to


"P_6(2,1,1,1)=\frac{6!}{2!1!1!1!}=\frac{720}2=360."

(b) The number of strings that can be formed by ordering the letters of the string CLASSICS is equal to


"P_8(3,2,1,1,1)=\frac{8!}{3!2!1!1!1!}=\frac{40320}{12}=3360."